OpenStudy (anonymous):

x/y + y/x = -1(x,y not equal to 0),then,the value of x^3-y^3 is?

OpenStudy (klimenkov):

\(x^3-y^3=(x-y)(x^2+xy+y^2)\)

OpenStudy (anonymous):

multiplying equation with xy we get. . . x^2+y^2=-xy x^2+y^2+xy=0 multiplying with (x-y) on both sides we get x^3+xy^2+x^2y-y^3-xy^2-x^y=0 that is x^3-y^3=0

OpenStudy (anonymous):

x^2+y^2+xy = 0 ??? how???????

OpenStudy (anonymous):

x^2+y^2=-xy x^2+y^2+xy=0 (taking -xy to left hand side

OpenStudy (anonymous):

\[\frac{x}{y} + \frac{y}{x} = -1 \implies \frac{x^2 + y^2}{xy} = -1 \implies x^2 + y^2 = -xy\] Add xy both sides: \[x^2 + y^2 + xy = 0\] Now use: \[\large x^3 - y^3 = (x-y)(x^2 + y^2 +xy) \implies \color{blue}{(x-y)(0)}\] multiply them you will have your answer..

OpenStudy (anonymous):

now i ge it thanks

OpenStudy (anonymous):

Welcome dear..